OpenStudy (anonymous):

Solve –3x^2 – 4x – 4 = 0

OpenStudy (anonymous):

(-3x+2)(x-2)=0 x=2/3 or x=2

OpenStudy (anonymous):

A concession stand charges $1.50 per soda and $2.00 per pretzel. Write an expression for the amount of money you would need to buy x sodas and y pretzels.

OpenStudy (anonymous):

1.50x + 2.00y